# Coarse-graining nonequilibrium diffusions with Markov chains

### v2

@@ 6. Discussion
In this paper, we presented a new approach to analytically derive a discrete-state Markov process from stationary planar diffusion, using finite-volume methods. Moreover, we showed that the EPR of the approximation converges in the infinite-state limit, which we illustrated with both solvable and unsolvable examples of stationary diffusions. In the case of unsolvable processes, this approximation gives an accurate way to investigate the behaviour of the EPR under changes to the parameters, which we illustrated with the stochastic van der Pol and Kuramoto models. Finally, we considered the problem of statistical inference from trajectory data. Using sampled paths, our results show that discrete-state models significantly underestimate the EPR of a NESS, even for long trajectories and fine grids. Nevertheless, discrete-state models can be used to define a hypothesis test that can be used to determine if a trajectory is from a NESS or not, which we, again, illustrated with sampled trajectories. Finally, we considered an illustrative application to real-world trajectories by analysing the group-polarisation trajectories from schooling fish [33].
@@
R.N.K completed research and wrote the manuscript. R.L. and A.G. supervised research and edited the manuscript.
@@ A.2. Solving for the stationary distribution
Solving for the stationary distribution involves solving the equation $L\pi=0$. Whilst for small number of discrete-states this can be calculated numerically in an efficient and exact way, for large state spaces (fine grids) this calculation becomes prohibitively slow. Instead, we solve
@@ B.1. Discretising non-diagonal diffusion with a nine-point stencil
(78) $\partial_{y}\left(D^{xy}p\right)|_{i+\frac{1}{2},j}$ $\approx\frac{D^{x}_{i+\frac{1}{2},j+1}p_{i+\frac{1}{2},j+1}-D^{x}_{i+\frac{1}{2},j-1}p_{i+\frac{1}{2},j-1}}{2\Delta x},$
(79) $\approx\frac{D^{x}_{i+\frac{1}{2},j+1}\left(p_{i+1,j+1}+p_{i,j+1}\right)-D^{x}_{i+\frac{1}{2},j-1}\left(p_{i+1,j-1}+p_{i,j-1}\right)}{4\Delta x}.$
@@ B.2. Diffusion which is diagonal under a coordinate transform
For a diffusion matrix which does not have constant eigenvectors, it would be best approximated with an adaptive, irregular grid. However, this severely complicated both theoretical and numerical analysis.

### v3

@@ 6. Discussion
In this paper, we presented a new approach to analytically derive a discrete-state Markov process from a stationary planar diffusion, using finite-volume methods. Moreover, we showed that the EPR of the approximation converges in the infinite-state limit, which we illustrated with both solvable and unsolvable examples of stationary diffusions. In the case of unsolvable processes, this approximation gives an accurate way to investigate the behaviour of the EPR under changes to the parameters, which we illustrated with the stochastic van der Pol and Kuramoto models. Finally, we considered the problem of statistical inference from trajectory data. Using sampled paths, our results show that discrete-state models significantly underestimate the EPR of a NESS, even for long trajectories and fine grids. Nevertheless, discrete-state models can be used to define a hypothesis test that can be used to determine if a trajectory is from a NESS or not, which we, again, illustrated with sampled trajectories. Finally, we considered an illustrative application to real-world trajectories by analysing the group-polarisation trajectories from schooling fish [33].
@@
R.N.K. completed research and wrote the manuscript. R.L. and A.G. supervised research and edited the manuscript.
@@ A.2. Solving for the stationary distribution
Solving for the stationary distribution involves solving the equation $L\pi=0$. Whilst for a small number of discrete-states this can be calculated numerically in an efficient and exact way, for large state spaces (fine grids) this calculation becomes prohibitively slow. Instead, we solve
@@ B.1. Discretising non-diagonal diffusion with a nine-point stencil
(78) $\partial_{y}\left(D^{xy}p\right)|_{i+\frac{1}{2},j}$ $\approx\frac{D^{xy}_{i+\frac{1}{2},j+1}p_{i+\frac{1}{2},j+1}-D^{xy}_{i+\frac{1}{2},j-1}p_{i+\frac{1}{2},j-1}}{2\Delta y},$
(79) $\approx\frac{D^{xy}_{i+\frac{1}{2},j+1}\left(p_{i+1,j+1}+p_{i,j+1}\right)-D^{xy}_{i+\frac{1}{2},j-1}\left(p_{i+1,j-1}+p_{i,j-1}\right)}{4\Delta y}.$
@@ B.2. Diffusion which is diagonal under a coordinate transform
For a diffusion matrix which does not have constant eigenvectors, it would be best approximated with an adaptive, irregular grid. However, this severely complicates both theoretical and numerical analysis.